# Nigel - Mechatronic Design and Robust Sim2Real Control of an Over-Actuated Autonomous Vehicle

### v1

@@ III Dynamics Modeling and Analysis
Dimensions (L×$\times$×B×$\times$×H) 0.31847$\times$0.175$\times$0.2568 m
@@
Friction coefficient (nominal) $\mu_{i}$ 0.4
[…]
We analyzed the model obtained in Eq. 2 based on the parameters identified (refer Table II) from lab experiments. The system has a stable equilibrium point under nominal conditions (ref Fig. 5(a)), however, drastic variation in friction or velocity significantly affects the system’s damping ratio, $\zeta$ (refer Fig. 5(b)). This follows that the system is prone to face sim2real transfer issues if the environmental conditions ($\mu_{i}$) are different from those in simulation or are uncertain in general, even at a constant velocity ($v$).
Figure 5: Analysis of vehicle dynamics model: (a) system phase portrait; and (b) system damping ratio as a function of friction and velocity.
@@ IV Robust Sim2Real Control Framework
Based on the linearized vehicle dynamics model derived in Section III a generalized representation of the open-loop system $S_{ol}$ is formulated (refer Eq. 3).
@@
Here, $\mathbf{x_{p}}=\begin{bmatrix}\beta&\dot{\psi}\end{bmatrix}^{T}$ are states, $\mathbf{u}=\begin{bmatrix}\delta_{FL}&\delta_{FR}&\delta_{RL}&\delta_{RR}\end{bmatrix}^{T}$ are control inputs and $\mathbf{w}=F_{w}$ is exogenous disturbance. The state transition is governed by $\mathbf{A_{p}}=\begin{bmatrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{bmatrix}$, $\mathbf{B_{p}}=\begin{bmatrix}b_{11}&b_{12}&b_{13}&b_{14}\\ b_{21}&b_{22}&b_{23}&b_{24}\end{bmatrix}$ and $\mathbf{D_{p}}=\begin{bmatrix}d_{1}\\ d_{2}\end{bmatrix}$. System outputs $\mathbf{y_{1}}=\mathbf{y_{2}}=\begin{bmatrix}\mathbf{x_{p}}&\mathbf{u}\end{bmatrix}^{T}$ are governed by $\mathbf{C_{p_{1}}}=\mathbf{C_{p_{2}}}=\begin{bmatrix}\mathbf{I}_{2\times 2}\\ \mathbf{0}_{4\times 2}\end{bmatrix}$, $\mathbf{B_{y_{1}}}=\mathbf{B_{y_{2}}}=\begin{bmatrix}\mathbf{0}_{2\times 4}\\ \mathbf{I}_{4\times 4}\end{bmatrix}$ and $\mathbf{D_{y}}=\mathbf{0}_{6\times 1}$. The measurements comprise full-state feedback with $\mathbf{M_{p}}=\mathbf{I}_{2\times 2}$ and $\mathbf{D_{z}}=\mathbf{0}_{2\times 1}$.
Considering the sim2real gap in terms of uncertainties in frictional coefficients of the 4 road-tire interconnects $\rho=\left<\mu_{FL},\mu_{FR},\mu_{RL},\mu_{RR}\right>$, where parameters $\rho_{i}$ can be time-varying or constant but uncertain, this work adopts polytopic modeling method for uncertainty treatment (refer Fig. 6). To this end, a polytopic linear parameter-varying system can be established (refer Eq. 4), where the state-space matrices depend affinely on the uncertain parameters.
@@
[…]
@@
Lemma 2: Considering the polytopic system presented in Eq. 6, the following statements are equivalent.
@@
    The system is stable with a quadratic $H_{\infty}$ performance index $\gamma_{1}$
@@
    $\Gamma_{ee}<\gamma_{1}$
  3. 3.
    There exists a single matrix $\mathbf{P}>0$ such that the following LMI is satisfied:
@@
  4. 4.
    There exists $\mathbf{P}>0$ such that the following system of LMIs is satisfied for $i=1,2,...,k$:
    $\begin{bmatrix}\mathbf{P\widetilde{A}}_{i}+\mathbf{\widetilde{A}}^{T}_{i}\mathbf{P}&\mathbf{P\widetilde{B}}_{i}&\mathbf{\widetilde{C}}^{T}_{i}\\ \ast&-\gamma_{1}\mathbf{I}&\mathbf{\widetilde{D}}^{T}_{i}\\ \ast&\ast&-\gamma_{1}\mathbf{I}\\ \end{bmatrix}<0$ (12)
@@
Lemma 3: Following statements are equivalent considering the polytopic system presented in Eq. 6.
@@
    The system is stable with a quadratic $H_{2}$ performance index $\gamma_{2}$
@@
    $\Gamma_{ep}<\gamma_{2}$
  3. 3.
    There exist $\mathbf{P}>0$ and $\mathbf{Q}>0$ such that the following conditions are satisfied:
@@
  4. 4.
    There exist $\mathbf{P}>0$ and $\mathbf{Q}>0$ such that the following sets of conditions are satisfied for $i=1,2,...,k$:
    $\begin{matrix}\mathrm{trace}\left[\mathbf{\widetilde{C}}_{i}\mathbf{P}\mathbf{\widetilde{C}}^{T}_{i}\right]<\gamma_{2}^{2}\\ \mathbf{\widetilde{A}}_{i}\mathbf{P}+\mathbf{P}\mathbf{\widetilde{A}}^{T}_{i}+\mathbf{\widetilde{B}}_{i}\mathbf{\widetilde{B}}^{T}_{i}<0\\ \mathrm{trace}\left[\mathbf{\widetilde{B}}^{T}_{i}\mathbf{Q}\mathbf{\widetilde{B}}_{i}\right]<\gamma_{2}^{2}\\ \mathbf{\widetilde{A}}^{T}_{i}\mathbf{Q}+\mathbf{Q}\mathbf{\widetilde{A}}_{i}+\mathbf{\widetilde{C}}^{T}_{i}\mathbf{\widetilde{C}}_{i}<0\end{matrix}$ (15)
@@
It can be shown that applying Eq. 16 to Eq. 11-12, we can arrive at Eq. 17a-17d and much in the same way, by applying Eq. 16 to Eq. 14-15, we can obtain Eq. 18a-18d.
Theorem 1: For a given plant of order $n_{p}$, there exists an $H_{\infty}$ controller of order $n_{c}\leqslant n_{p}$ to stabilize the closed-loop system and guarantees $\Gamma_{ee}<\gamma_{1}$ if and only if the conditions presented in Eq. 17a-17d are satisfied.
@@
Remark 1: The constraints imposed by Eq. 17a-17c are convex, whereas Eq. 17d being rank constraint, is not convex.
Remark 2: For a full-order controller, i.e., $n_{c}=n_{p}$, constraint imposed by Eq. 17d is satisfied and the resultant overall problem becomes convex.
Remark 3: For full-state feedback controller, i.e., $n_{c}=0$ and $\mathbf{M_{p}}=\mathbf{I}$ constraints imposed by Eq. 17b-17d are satisfied and the resultant overall problem becomes convex.
Theorem 2: For a given plant of order $n_{p}$, there exists an $H_{2}$ controller of order $n_{c}\leqslant n_{p}$ to stabilize the closed-loop system and guarantees $\Gamma_{ep}<\gamma_{2}$ if and only if the conditions presented in Eq. 18a-18d are satisfied.
@@
Remark 4: The constraints imposed by Eq. 18a-18c are convex, whereas Eq. 18d being rank constraint, is not convex.
Remark 5: For a full-order controller, i.e., $n_{c}=n_{p}$, constraint imposed by Eq. 18d is satisfied and the resultant overall problem becomes convex.
Remark 6: For full-state feedback controller, i.e., $n_{c}=0$ and $\mathbf{M_{p}}=\mathbf{I}$ constraints imposed by Eq. 18b-18d are satisfied and the resultant overall problem becomes convex.
From the solution $\left\{\mathbf{X},\mathbf{Y}\right\}$ to the convex optimization problem of minimizing $\mathrm{trace}\left(\mathbf{X}+\mathbf{Y}\right)$ subject to the inequalities presented in Eq. 17a and Eq. 18a, we can build the Lyapunov matrix and obtain the solution for the optimal state-feedback controller $\mathbf{K}$ whose purpose is to minimize the influence of the disturbance $\mathbf{w}$ on the system response $\mathbf{y}=\begin{bmatrix}\mathbf{y_{1}}&\mathbf{y_{2}}\end{bmatrix}^{T}$ resulting in a closed-loop system $S_{cl}$ with $\mathbf{A_{cl}}=\left(\mathbf{A_{p}}+\mathbf{B_{p}}\mathbf{K}\right)$. That is to say that we obtain an optimal state-feedback control $\mathbf{u}=\mathbf{Kx_{p}}$ that:
@@
Remark 7: With the specific choice of output quantities $\mathbf{y}=\begin{bmatrix}\mathbf{y_{1}}&\mathbf{y_{2}}\end{bmatrix}^{T}$ as stated earlier, we can see that the designed controller penalizes deviation of state $\mathbf{x_{p}}$ from zero (for stabilization) and reference $\mathbf{x_{r}}$ (for tracking) along with the 4 control inputs $\mathbf{u}=\begin{bmatrix}\delta_{FL}&\delta_{FR}&\delta_{RL}&\delta_{RR}\end{bmatrix}^{T}$ so as to conserve energy by reducing the control effort.
@@ V Results and Discussion
Figure 7: The simulation experiments comprised testing the controller for (a) stabilizing as well as steady-state tracking of standard benchmark maneuvers viz. (b) lane-change, (c) skidpad, (d) fishhook and (e) slalom tests, against varying friction for individual road-tire interconnects and wind-gust disturbances.
Figure 8: The real-world experiment comprised testing the controller for steady-state tracking of a “Figure-8” maneuver, against varying friction for individual road-tire interconnects and poking disturbances: (a) depicts vehicle states and trajectory; and (b) visualizes the vehicle trajectory as a time-series snapshot, with the inset indicating the disturbance $F_{w}$ and soap-water patches of different concentrations to vary road friction $\mu$.
The closed-loop system derived in Section IV was found to be robustly stable over the entire specified parameter range, since a single Lyapunov function of the form $\mathbf{V}(\mathbf{x_{p}},\alpha)=\mathbf{x_{p}}^{T}\mathbf{Q}^{-1}(\alpha)\mathbf{x_{p}},\>\mathbf{Q}(\alpha)=\alpha_{1}\mathbf{Q_{1}}+...+\alpha_{k}\mathbf{Q_{k}}$ was obtained such that $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{V}(\mathbf{x_{p}},\alpha)<0$ for all polytopic decompositions.
@@
We analyzed the performance of the robust optimal controller in simulation as well as real-world settings. The design of experiments followed a similar approach for both. The stabilizing controller was tested for a straight-line maneuver. For the tracking controller, reference generation was achieved by running the open-loop system with standard test signals viz. impulse, step, ramp and sine, which resulted in benchmark maneuvers viz. lane-change, skidpad, fishhook and slalom, respectively in the absence of any disturbances or parameter variations. It is worth mentioning that the controllers discussed in this work are designed to track the vehicle states, which in turn results in spatiotemporal trajectory tracking.
The simulation experiments comprised testing the controller for stabilizing as well as steady-state tracking of standard benchmark maneuvers viz. lane-change, skidpad, fishhook and slalom, against environmental uncertainties (refer Fig. 7). Particularly, the friction for individual road-tire interconnects was simulated as phase-shifted sinusoids, such that all 4 wheels never experienced the same $\mu$, and wind-gust disturbances were simulated as delayed step input between $t$=7.5 to $t$=22.5 seconds. Apart from the results presented in this section, the controller was tested exhaustively for varying velocities of the vehicle, with values as high as 10 times the nominal operating limit. Additionally, we also synthesized and validated a similar robust optimal controller for nominal full-scale vehicle parameters. Analysis revealed that the same design framework (refer Fig. 6) is effective for such extensions.
@@
Upon validating the controller performance in simulation, we deployed the controller on the physical prototype of Nigel and validated its performance in a real-world experiment. The said experiment involved steady-state tracking of “Figure-8” maneuver, where the state feedbacks were obtained using the OptiTrack motion capture system (refer Fig. 8). In order to widen the sim2real gap, the test surface was deliberately left unclean with dust/sand particles, irregular scratches and tape residues from previous experiments. Additionally, soap-water solutions of varying concentrations were spilled in uneven quantities on the anticipated path, to induce uncertainty in terms of road friction. Finally, similar to the simulation experiments, the vehicle was disturbed actively by poking it with a pole at $t$=52, $t$=68, $t$=81 and $t$=109 seconds, which acted as an impulse input.

### v3

@@ III Dynamics Modeling and Analysis
Dimensions (L×$\times$×B×$\times$×H) 0.318$\times$0.175$\times$0.257 m
@@
Friction coefficient $\mu_{i}$ 0.4 (nominal)
Tire stiffness $C_{i}$ 22.4768 N.rad-1
We analyzed the model obtained in Eq. 2 based on the parameters identified (refer Table II) from lab experiments. The system has a stable equilibrium point under nominal conditions (refer Fig. 5(a)), but drastic variation in operating conditions (friction or velocity) significantly affects the system’s damping ratio, $\zeta$ = $\frac{-\lambda_{R}}{\sqrt{\lambda_{R}^{2}+\lambda_{I}^{2}}}$ (refer Fig. 5(b)) where $\lambda=\lambda_{R}\pm\lambda_{I}\iota$ are eigenvalues of the system matrix. This follows that the system is prone to face sim2real transfer issues if the real-world conditions ($\mu_{i}$) are different from the simulation or uncertain in general, even at a constant velocity ($v$).
Figure 5: Analysis of vehicle dynamics: (a) system phase portrait @ $\mu$ = 0.4, $v$ = 0.35 m/s; and (b) system damping ratio $\zeta$ as a function of $\mu$ and $v$.
@@ IV Robust Sim2Real Control Framework
The proposed framework (see Fig. 6) begins with formulating a linear parameter-varying (LPV) description of the system in polytopic form. Then, a robust optimal controller is synthesized by solving a semi-definite program using linear matrix inequalities (LMIs). Finally, the closed-loop system response is verified in simulation, followed by sim2real validation.
To this end, a generalized representation of the open-loop system $S_{ol}$ is obtained (refer Eq. 3).
@@
Here, $\mathbf{u}=\begin{bmatrix}\delta_{FL}&\delta_{FR}&\delta_{RL}&\delta_{RR}\end{bmatrix}^{T}$ are the control inputs, $\mathbf{x_{p}}=\begin{bmatrix}\beta&\dot{\psi}\end{bmatrix}^{T}$ are the states and $\mathbf{w}=F_{w}$ is the disturbance.
The state transition is governed by $\mathbf{A_{p}}=\begin{bmatrix}a_{11}&a_{12}\\ a_{21}&a_{22}\end{bmatrix}$, $\mathbf{B_{p}}=\begin{bmatrix}b_{11}&b_{12}&b_{13}&b_{14}\\ b_{21}&b_{22}&b_{23}&b_{24}\end{bmatrix}$ and $\mathbf{D_{p}}=\begin{bmatrix}d_{1}\\ d_{2}\end{bmatrix}$. System outputs $\mathbf{y_{1}}=\mathbf{y_{2}}=\begin{bmatrix}\mathbf{x_{p}}&\mathbf{u}\end{bmatrix}^{T}$ are governed by $\mathbf{C_{p_{1}}}=\mathbf{C_{p_{2}}}=\begin{bmatrix}\mathbf{I}_{2\times 2}\\ \mathbf{0}_{4\times 2}\end{bmatrix}$, $\mathbf{B_{y_{1}}}=\mathbf{B_{y_{2}}}=\begin{bmatrix}\mathbf{0}_{2\times 4}\\ \mathbf{I}_{4\times 4}\end{bmatrix}$ and $\mathbf{D_{y}}=\mathbf{0}_{6\times 1}$. The measurements comprise full-state feedback with $\mathbf{M_{p}}=\mathbf{I}_{2\times 2}$ and $\mathbf{D_{z}}=\mathbf{0}_{2\times 1}$.
Considering the sim2real gap in terms of uncertainties in frictional coefficients of the 4 road-tire interconnects $\rho=\left<\mu_{FL},\mu_{FR},\mu_{RL},\mu_{RR}\right>$, where parameters $\rho_{i}$ can be time-varying or constant but uncertain, this work adopts polytopic modeling method for uncertainty treatment. To this end, a polytopic LPV system can be established (refer Eq. 4), where the state-space matrices depend affinely on the uncertain parameters.
@@
We intend to place poles of the closed-loop system $S_{cl}$ in the LMI region governed by the intersection of left half-plane with $\Re(z)<\alpha$ (i.e., $\alpha$-stability), where $\alpha=-0.1$, and a conic sector centered at the origin having an inner angle of $\phi=3\pi/4$ so as to guarantee some minimum decay rate and closed-loop damping. The LMI in Eq. 9a represents the $\alpha$-stability region, while the one in Eq. 9b represents the conic sector centered at the origin with an inner angle of $\phi=2\theta$.
@@
Lemma 2: Considering the polytopic system described by Eq. 6, the following statements are equivalent.
@@
    The system is stable with a quadratic $H_{\infty}$ performance index $\gamma_{1}$.
@@
    There exists a matrix $\mathbf{P}>0$ such that the following LMI has a feasible solution:
@@
Lemma 3: Following statements are equivalent for the uncertain system described by the polytopic model in Eq. 6.
@@
    The system is stable with quadratic $H_{2}$ performance $\gamma_{2}$.
@@
    There exist $\mathbf{P}>0$ and $\mathbf{Q}>0$ such that:
@@
It can be shown that applying Eq. 14 to Eq. 11, we can arrive at Eq. 15a-15d and much in the same way, by applying Eq. 14 to Eq. 13, we can obtain Eq. 16a-16d.
Theorem 1: For a given plant of order $n_{p}$, there exists an $H_{\infty}$ controller of order $n_{c}\leqslant n_{p}$ to stabilize the closed-loop system and guarantees $\Gamma_{ee}<\gamma_{1}$ if and only if the conditions presented in Eq. 15a-15d are satisfied.
@@
Remark 1: For the full-state feedback controller, i.e., $n_{c}=0$ and $\mathbf{M_{p}}=\mathbf{I}$, constraints imposed by Eq. 15b-15d are satisfied, and the resultant overall problem becomes convex.
Theorem 2: For a given plant of order $n_{p}$, there exists an $H_{2}$ controller of order $n_{c}\leqslant n_{p}$ to stabilize the closed-loop system and guarantees $\Gamma_{ep}<\gamma_{2}$ if and only if the conditions presented in Eq. 16a-16d are satisfied.
@@
Remark 2: For the full-state feedback controller, i.e., $n_{c}=0$ and $\mathbf{M_{p}}=\mathbf{I}$, constraints imposed by Eq. 16b-16d are satisfied, and the resultant overall problem becomes convex.
From the solution $\left\{\mathbf{X},\mathbf{Y}\right\}$ to the convex optimization problem of minimizing $\mathrm{trace}\left(\mathbf{X}+\mathbf{Y}\right)$ subject to Eq. 15a and Eq. 16a, we can build the Lyapunov matrix and obtain the optimal state-feedback controller $\mathbf{K}$ whose purpose is to minimize the influence of disturbance $\mathbf{w}$ on the uncertain system response $\mathbf{y}=\begin{bmatrix}\mathbf{y_{1}}&\mathbf{y_{2}}\end{bmatrix}^{T}$ resulting in a closed-loop system $S_{cl}$ with $\mathbf{A_{cl}}=\left(\mathbf{A_{p}}+\mathbf{B_{p}}\mathbf{K}\right)$. In other words, we obtain an optimal feedback matrix $\mathbf{K}$ that achieves the following.
@@
Remark 3: With the specific choice of output signals $\mathbf{y}=\begin{bmatrix}\mathbf{y_{1}}&\mathbf{y_{2}}\end{bmatrix}^{T}$, the designed controller minimizes the deviation of state $\mathbf{x_{p}}$ from zero (for stabilization) or reference $\mathbf{x_{r}}$ (for tracking) along with the control inputs $\mathbf{u}$. From a sim2real transfer perspective, the devised framework guarantees stabilizing/tracking performance in the presence of parameter uncertainties and disturbances while also minimizing the control effort.
@@ V Results and Discussion
We first established the benefit of adopting an unconventional vehicle architecture by comparing it against a conventional one with the same parameters (refer Table II). The conventional Ackermann-steered architecture yielded a Pareto-optimal $H_{\infty}$ performance of 2.61e-1 and $H_{2}$ performance of 6.45e-1, while the unconventional independent 4WD4WS architecture was able to guarantee $H_{\infty}$ performance of 1.98e-01 and $H_{2}$ performance of 5.56e-01. In other words, the conventional system was poorly robust as compared to the proposed one and was, therefore, capable of handling relatively narrower sim2real gaps. Consequently, further results are reported only for the independent 4WD4WS vehicle architecture.
TABLE III: Benchmarking analysis of the proposed sim2real framework
Error ($\varepsilon$) $\searrow$ Open-Loop Non-Robust Robust
Sim Real Sim Real Sim Real
Straight 3.81e-2 5.03e-2 9.73e-2 6.04e-2 1.83e-2 1.82e-2
Lane-Change 7.59e-2 6.59e-2 5.78e-2 3.38e-2 1.20e-2 1.18e-2
Skidpad 1.09e-1 1.01e-1 7.91e-2 8.36e-2 2.14e-2 2.06e-2
Fishhook 5.52e-2 6.79e-2 4.23e-2 7.75e-2 1.50e-2 1.82e-2
Slalom 1.82e-1 1.91e-1 2.87e-1 2.52e-1 6.14e-2 6.15e-2
Figure-8 1.43e-1 4.15e-1 6.85e-2 5.47e-1 4.67e-2 1.07e-1
Next, we benchmarked the performance of the proposed robust sim2real control framework against (i) replayed reference control commands in open-loop and (ii) another non-robust state-feedback controller devised with a pole-placement approach based on Eq. 9a to ensure $\Re(z)<-2.0$. Particularly, we analyzed the root-mean-squared error (RMSE), $\epsilon=\begin{bmatrix}\epsilon_{x}\;\epsilon_{y}\;\epsilon_{\psi}\\ \end{bmatrix}^{T}$, for the Cartesian pose of the vehicle throughout a given trajectory. The single error metric reported in Table III is the $\ell^{2}$-norm of the resulting errors in each of the pose coordinates, i.e., $\varepsilon=\left\|\epsilon\right\|_{2}$.
@@
The design of experiments followed a common approach for simulation and real-world settings. The stabilizing controller was tested for a straight-line maneuver (refer Fig. 7(a)). The steady-state tracking control was validated for standard benchmark maneuvers, viz. lane-change (refer Fig. 7(b)), skidpad (refer Fig. 7(c)), fishhook (refer Fig. 7(d)), slalom (refer Fig. 7(e)) and figure-8 (refer Fig. 7(f)), wherein the reference generation was achieved by simulating the open-loop system with standard test signals, viz. impulse, step, ramp and sine.
Figure 7: Experimentation and benchmarking results from simulation and real-world for (a) stabilizing as well as (b)-(f) steady-state tracking of standard maneuvers, viz. (b) lane-change, (c) skidpad, (d) fishhook, (e) slalom tests, and (f) figure-8, under deliberately injected exogenous disturbances and uncertainties.
In simulation experiments, the friction for individual road-tire interconnects was exaggerated as phase-shifted sinusoids with $\mu_{i}=0.35\sin(2\pi f\theta-\phi)+0.55\in\left[0.2,0.9\right]$, where $f$=12 is the frequency of variation, $\theta$=$[0,2\pi]$ is the range of variation and $\phi$ is the phase lag. The resulting sinusoids were injected with synthetic noise defined as $\mu_{i}-0.05+0.1\chi$, where $\chi\sim U(0,1)$ is a uniform random variable. Here, $\mu_{FR}$, $\mu_{RL}$ and $\mu_{RR}$ respectively lagged $\mu_{FL}$ by $\phi_{FR}=\pi/2$, $\phi_{RL}=\pi$ and $\phi_{RR}=3\pi/2$ rad, such that all four wheels never experienced the same $\mu$. Additionally, noisy wind-gust disturbance $F_{w}$ of $0.25$ N amplitude was simulated as a step input from $t$=1 second onwards, with synthetic noise of the form $F_{w}+0.1\chi*\mathrm{max}(F_{w})$, where $\chi\sim U(0,1)$ is a uniform random variable.
@@
In real-world experiments, the vehicle state feedback was obtained using OptiTrack motion capture system. Although a nominal value of friction coefficient was estimated to be 0.4 from the motion capture data (refer Table II), the test surface (5.0$\times$3.5 m) was deliberately left unclean with dust/sand particles, irregular scratches and tape residues from previous experiments. Additionally, soap-water solutions of varying concentrations were spilled in uneven quantities to aggravate the uncertainty in terms of road friction. Finally, in the figure-8 maneuver (depicted as a freeze-frame sequence in Fig. 1), the vehicle was actively poked with a pole at $t$=52, $t$=68, $t$=81 and $t$=109 seconds, which acted as an impulse disturbance. The exact values of real-world friction and disturbance were time-varying and very difficult to measure and were therefore treated to be a part of the “unknown” sim2real gap.
Quantitatively, the proposed robust sim2real framework has the least error across all maneuvers for all experiments. Additionally, the difference between “Sim” and “Real” error metrics practically reflects the sim2real gap, wherein the proposed framework clearly outperforms its counterparts.
Qualitatively, all the systems track the first lane-change transient relatively closely, but the open-loop and non-robust systems digress significantly after the second one. For skidpad and fishhook maneuvers, all the system performances overlap before and within the loops, albeit with slight variations, after which we can see a drastic variation, as different systems exit at different angles. For straight and slalom maneuvers, the open-loop system deceptively seems to outperform the non-robust closed-loop system, but note that it is not actively correcting any errors. Finally, figure-8 maneuver was the only one where the real-world system was actively disturbed and, consequently, the simulated and real-world performance have a wider gap in this case. Nevertheless, the proposed framework still outperforms the open-loop and non-robust systems.